# REPLAY: Modeling Time-Varying Temporal Regularities of Human Mobility for Location Prediction over Sparse Trajectories

### v1

@@ 3.1 Smoothed Timestamp Embeddings
Hour-in-Week Timestamp Transformation. For $i$-th check-in in the input mobility trace, we first transform its check-in time $t_{i}$ into an hour-in-week timestamp $t_{n}$, where $1\leq n\leq 168$ (i.e., one of the 168 hours in a week). For example, $t_{i}=$ (15:00 Tue. 07/11/2023) is transformed into $t_{n}$, $n=39$, as $t_{i}$ is the 39-th hour in a week. We choose to model the temporal regularity at an hour granularity on a weekly scale, as it can capture not only daily patterns but also the difference between weekdays and weekends, which is also suggested by Yang et al. (2019). Moreover, we experimentally justify this design choice by comparing with other time granularities and scales (see experiments and Appendix D for more detail).
Gaussian Smoothing with Learnable Bandwidth. For each timestamp, we initialize a learnable embedding vector, denoted as $\vec{t}_{n}$, and compute a smoothed timestamp embedding $\vec{s}_{n}$ using Gaussian weighted averaging with a learnable bandwidth $\sigma_{n}$, flexibly combining the neighboring timestamp embeddings. This design is inspired by kernel regression techniques Nadaraya (1964), and we further incorporate timestamp-specific learnable bandwidths here to flexibly capture the temporal regularities of different strengths across different timestamps. Specifically, for the input timestamp $t_{n}$, the Gaussian kernel function is defined as:
@@
Cyclical Timestamps Different from traditional distance metrics in kernel regression, our distance function for hour-in-week timestamps should be carefully designed, as Sunday night 11pm ($l=168$) and Monday morning 1am ($l=2$) are only two hours apart, for example. Therefore, we consider cyclical timestamps for the distance computation:
@@
3.2 Flashback with Smoothed Timestamp Embeddings
Flashback Mechanism After obtaining the smoothed timestamp embedding, we feed it together with the corresponding POI embedding to RNN units with the flashback mechanism. Specifically, the flashback mechanism Yang et al. (2020a) leverages the spatiotemporal context to search past hidden states with high predictive power. It computes the weighted average of the historical hidden states $h_{j}$, $j<i$, with a weight $\mathcal{W}(\Delta T_{i,j},\Delta D_{i,j})$ as an aggregated hidden state for location prediction. This weight is parameterized by the temporal and spatial distances ($\Delta T_{i,j}$ and $\Delta D_{i,j}$, respectively) between check-in $(p_{i},t_{i})$ and $(p_{j},t_{j})$, which is defined as follows:
@@ 3.2 Flashback with Smoothed Timestamp Embeddings
where $\mbox{hvc}(x)=\frac{1+\cos(x)}{2}$ is the Havercosine function capturing the daily periodicity (note that $T_{i,j}$ is measured in days here). The two exponential terms model the temporal and spatial decay of past hidden states, respectively, following the intuition that the hidden states of older and farther check-ins have less predictive power in general; $\alpha$ and $\beta$ are tunable parameters for decay rates. Please refer to Yang et al. (2020a) for more details.
Adding Smoothed Timestamp as Input Different from the original Flashback network Yang et al. (2020a) that only takes POI embeddings as inputs of RNN units, REPLAY concatenates the POI embedding together with the corresponding smoothed timestamp embedding as inputs of the RNN units, i.e., $\vec{h}_{i}=\mathcal{F}([\vec{p}_{i};\vec{s}_{n}],\vec{h}_{i-1})$, as shown in Figure 2. The objective here is to capture the correlation between a check-in’s POI and timestamp. Such a correlation can effectively help the location prediction performance. For example, working-related POIs are usually checked by users in the daytime. Moreover, our smoothed timestamp embeddings here can flexibly capture this correlation by integrating information from neighboring timestamps, where the amount of the information is controlled by the bandwidth $\sigma$. Subsequently, the learnable $\sigma$ can automatically adapt to time-varying temporal regularities; for example, a timestamp with a high mobility regularity will intuitively yield a smaller value of the bandwidth $\sigma$, as it relies less on the information from its neighboring timestamps for prediction.
@@ 3.3 Prediction Conditioned on Query Timestamp
The design choice of prediction conditioned on timestamp is motivated by the fact that over a sparse user mobility trace, the time intervals between successive check-ins significantly vary. Subsequently, to predict the next check-in in the sequence, the predicted next POI should vary across different “query” time. Therefore, REPLAY is designed to make predictions conditioned on a query time. The same or similar settings exist in the literature by using a query timestamp (same as ours) Gao et al. (2012); Yang et al. (2019); Li et al. (2021a), a query time interval (to predict the location of a user after a given time period) Figueiredo et al. (2016); Zhao et al. (2019, 2020); Feng et al. (2018), or a fixed query time threshold (to predict the location of a user within a forthcoming period of time) Feng et al. (2015); Zhang et al. (2014a); Feng et al. (2018).
Note that an alternative setting is to make predictions on the next POI without considering the query time Liu et al. (2016); Yang et al. (2020a); Rao et al. (2022). We do not advocate for this setting in this paper as it always makes the same prediction for all future time given an input mobility trace, which is less reasonable in real-world scenarios.
@@ 3.4 Discussions
Time interval captures the temporal regularity of human mobility over sparse trajectories. The flashback mechanism leverages the spatiotemporal context to search the historical hidden states for location prediction, where the historical hidden states temporally close to the periodicity peaks of returning probability (as shown in Figure 1) are usually more informative for location prediction. For example, Figure 3 shows that using time interval ($\Delta T\approx 1$ day), we can quickly find similar temporal context in sparse trajectories and retrieves the corresponding hidden state for prediction.
@@
Figure 3: A toy example showing how REPLAY models the time-varying temporal regularities of human mobility.
@@
Smoothed Timestamps further captures the temporal variation (uncertainty) of the temporal regularities. Specifically, we reveal in this paper that the degree of such uncertainty (or the strength of the temporal regularities) varies across time. Subsequently, the smoothed timestamp embeddings are used to accommodate such variations. Instead of directly using a specific timestamp embedding, we smooth it with the neighboring timestamps as shown in Figure 3, where the timestamp-specific learnable bandwidths can automatically adapt to the temporal regularities of different strengths across different timestamps.
Model Training. Please refer to Appendix A for more details about the model training.
Table 1: Location prediction performance comparison, where the best-performing ones are highlighted. ${}^{\dagger}$ethods make predictions using query time information as a query timestamp (MFT, LBSN2Vec, and REPLAY), as a query time interval (TribeFlow, DeepMove, STGN, STGCN, STAN), or as a query time threshold (PRME). Note that DeepMove predicts locations in a fixed future time slot, implicitly implying a fixed query time interval. ${}^{\diamond}$xperiments of GetNext are conducted on 15000 randomly sampled users on Foursquare, due to its poor efficiency where it takes more than 24 hours per epoch on the whole Foursquare datasets. The top two best-performing results are denoted in bold and underlined, respectively.
Method Gowalla Foursquare
Acc@1 Acc@5 Acc@10 MRR Acc@1 Acc@5 Acc@10 MRR
BPR 0.0131 0.0363 0.0539 0.0235 0.0315 0.0828 0.1143 0.0538
Feature-based Methods ${}^{\dagger}$ 0.0525 0.0948 0.1052 0.0717 0.1945 0.2692 0.2788 0.2285
${}^{\dagger}$ 0.0864 0.1186 0.1390 0.1032 0.2190 0.3955 0.4621 0.2781
${}^{\dagger}$ 0.0740 0.2146 0.2899 0.1504 0.0982 0.3167 0.4064 0.2040
${}^{\dagger}$ 0.0256 0.0723 0.1143 0.0583 0.0297 0.0832 0.1239 0.0645
LSTM 0.0621 0.1637 0.2182 0.1144 0.1144 0.2949 0.3761 0.2018
GRU 0.0528 0.1416 0.1915 0.0993 0.0606 0.1797 0.2574 0.1245
${}^{\dagger}$ 0.0625 0.1304 0.1594 0.0982 0.2400 0.4319 0.4742 0.3270
${}^{\dagger}$ 0.0624 0.1586 0.2104 0.1125 0.2094 0.4734 0.5470 0.3283
${}^{\dagger}$ 0.0546 0.1440 0.1932 0.1017 0.1878 0.4502 0.5329 0.3062
${}^{\dagger}$ 0.0891 0.2096 0.2763 0.1523 0.2265 0.4515 0.5310 0.3420
${}^{\diamond}$ 0.0912 0.2003 0.2487 0.1484 0.1862 0.4702 0.5763 0.3153
Flashback 0.1158 0.2754 0.3479 0.1925 0.2496 0.5399 0.6236 0.3805
${}^{\dagger}$ 0.1866 0.3476 0.4124 0.2635 0.3529 0.5953 0.6648 0.4638
@@ 4.1 Experiment Setup
Dataset. We adopt two widely used check-in datasets collected from two LBSNs: Gowalla Cho et al. (2011) and Foursquare Yang et al. (2020b). Table 2 shows the dataset statistics. Each mobility trace is chronologically split into 80% for training and 20% for testing. Note that as the user mobility regularity depends on local time, we transform the original UTC time of each check-in to a local time according to the time zone of the check-in GPS coordinates.
Table 2: Statistics of the datasets
@@
Baselines. We compare REPLAY against a sizable collection of 18 state-of-the-art techniques from five categories: User Preference-based Methods WRMF Hu et al. (2008) and BPR Rendle et al. (2009); Feature-based methods Most Frequent Time (MFT) Gao et al. (2012) and LBSN2Vec Yang et al. (2019); Markov-Chain-based Methods FPMC Rendle et al. (2010), PRME Feng et al. (2015), and TribeFlow Figueiredo et al. (2016); Basic RNNs: RNN Zhang et al. (2014b), LSTM (Long Short-Term Memory) Hochreiter and Schmidhuber (1997) and GRU (Gated Recurrent Unit) Cho et al. (2014); Spatiotemporal Sequence Models STRNN Liu et al. (2016), DeepMove Feng et al. (2018), STGN Zhao et al. (2019), STGCN Zhao et al. (2019), STAN Luo et al. (2021), GetNext Yang et al. (2022), Flashback Yang et al. (2020a), and Graph-Flashback Rao et al. (2022). The details about these baselines and their hyperparameter setting can be found in Appendix B.
Evaluation Metrics. We report two widely used metrics for location prediction: average Accuracy@N (Acc@N), where $N=1,5,10$, and Mean Reciprocal Rank (MRR). Please refer to Appendix C for more details about experiment settings.

### v3

@@ 3.1 Smoothed Timestamp Embeddings
3.1.1 Hour-in-Week Timestamp Transformation
For $i$-th check-in in the input mobility trace, we first transform its check-in time $t_{i}$ into an hour-in-week timestamp $t_{n}$, where $1\leq n\leq 168$ (i.e., one of the 168 hours in a week). For example, $t_{i}=$ (15:00 Tue. 07/11/2023) is transformed into $t_{n}$, $n=39$, as $t_{i}$ is the 39-th hour in a week. We choose to model the temporal regularity at an hour granularity on a weekly scale, as it can capture not only daily patterns but also the difference between weekdays and weekends, which is also suggested by [32]. Moreover, we experimentally justify this design choice by comparing it with other time granularities and scales (see Section 4.4 for more detail).
3.1.2 Gaussian Smoothing with Learnable Bandwidth
For each hour-in-week timestamp, we initialize a learnable embedding vector, denoted as $\vec{t}_{n}$, and then compute a smoothed timestamp embedding $\vec{s}_{n}$ using Gaussian weighted averaging with a learnable bandwidth $\sigma_{n}$, flexibly combining the neighboring timestamp embeddings. This design is inspired by kernel regression techniques [56], and we further incorporate timestamp-specific learnable bandwidths here to flexibly capture the temporal regularities of different strengths across different timestamps. Specifically, for the input timestamp $t_{n}$, the Gaussian kernel function is defined as:
@@ 3.1.2 Gaussian Smoothing with Learnable Bandwidth
3.1.3 Cyclical Timestamps
Different from traditional distance metrics in kernel regression, our distance function for hour-in-week timestamps should be carefully designed, as Sunday night 11pm ($l=168$) and Monday morning 1am ($l=2$) are only two hours apart, for example. Therefore, we consider cyclical timestamps for the distance computation:
@@ 3.1.3 Cyclical Timestamps
3.2 Flashback Network with Smoothed Timestamp Embeddings
After obtaining the smoothed timestamp embedding, we feed it together with the corresponding POI embedding to RNN units with the flashback mechanism. In the following, we first briefly present the Flashback mechanism, followed by our extension of its input to add smoothed timestamp embeddings.
3.2.1 Flashback Mechanism
Flashback [14] is a general RNN architecture leveraging the spatiotemporal information in sparse trajectories to search the informative past hidden states for location prediction. Specifically, it takes a sequence of check-in POIs $\{...,p_{i-2},p_{i-1},p_{i},...\}$ as the input of an RNN (vanilla RNN, GRU, or LSTM) and output its recurrent hidden state $\vec{h}_{i}$. Then, the flashback mechanism leverages the spatiotemporal distance information to search the informative past hidden states. To this end, we use learnable weight $\mathcal{W}(\Delta T_{i,j},\Delta D_{i,j})$ to re-weighting the historical hidden states $h_{j}(j<i)$ for location prediction. The weight used for aggregating the hidden state is parameterized by the spatial distance $\Delta D_{i,j}$ and temporal distance $\Delta T_{i,j}$ between the check-ins $(p_{i},t_{i})$ and $(p_{j},t_{j})$, which is defined as follows:
@@ 3.2.1 Flashback Mechanism
where the Havercosine function $\mbox{hvc}(x)=\frac{1+\cos(x)}{2}$ is used to capture the temporal periodicity ($\Delta T_{i,j}$ is measured in days here). This Havercosine function effectively models the daily periodicity of human mobility based on the temporal distances $\Delta T_{i,j}$ as shown in Figure 1. Note that the distance between transformed timestamps used in smoothed timestamp embeddings differs from the temporal distance in Flashback here; the former is used for smoothing timestamp embeddings to accommodate the time-varying temporal regularities, while the latter is used for weighing the past hidden state to capture the temporal periodicity. The two exponential terms model the temporal and spatial decay of past hidden states, respectively, following the intuition that the hidden states of older and farther check-ins have less predictive power in general; $\alpha$ and $\beta$ are tunable parameters for decay rates. Please refer to [14] for more details.
3.2.2 Adding Smoothed Timestamp as Input
Different from the original Flashback network [14] that only takes POI embeddings as inputs of RNN units, REPLAY concatenates the POI embedding together with the corresponding smoothed timestamp embedding as inputs of the RNN units, i.e., $\vec{h}_{i}=\mathcal{F}([\vec{p}_{i};\vec{s}_{n}],\vec{h}_{i-1})$, as shown in Figure 2. The objective here is to capture the correlation between a check-in’s POI and timestamp. Such a correlation can effectively help the location prediction performance. For example, working-related POIs are usually checked by users in the daytime. Moreover, our smoothed timestamp embeddings here can flexibly capture this correlation by integrating information from neighboring timestamps, where the amount of the information is controlled by the bandwidth $\sigma$. Subsequently, the learnable $\sigma$ can automatically adapt to time-varying temporal regularities; for example, a timestamp with a high mobility regularity will intuitively yield a smaller value of the bandwidth $\sigma$, as it relies less on the information from its neighboring timestamps for prediction.
@@ 3.3 Prediction Conditioned on Query Timestamp
The design choice of prediction conditioned on timestamp is motivated by the fact that over a sparse user mobility trace, the time intervals between successive check-ins significantly vary. Subsequently, to predict the next check-in in the sequence, the predicted next POI should vary across different “query” time. Therefore, REPLAY is designed to make predictions conditioned on a query time. The same or similar settings exist in the literature by using a query timestamp (same as ours) [37, 32, 15], a query time interval (to predict the location of a user after a given time period) [57, 12, 13, 58], or a fixed query time threshold (to predict the location of a user within a forthcoming period of time) [49, 59, 58].
Note that an alternative setting is to make predictions on the next POI without considering the query time [7, 14, 17]. We do not advocate for this setting in this paper as it always makes the same prediction for all future time given an input mobility trace, which is less reasonable in real-world scenarios.
@@ 3.4 Discussions
Time interval captures the temporal regularity of human mobility over sparse trajectories. Real-world mobility trajectories are often sparse and incomplete, which hinders the application of recurrent neural network models for mobility modeling. Our flashback mechanism is able to effectively utilize the rich spatiotemporal context to search the historical hidden states for location prediction, where the historical hidden states temporally close to the periodicity peaks of returning probability (as shown in Figure 1) are usually more informative for location prediction. For example, Figure 3 shows that using time interval ($\Delta T\approx 1$ day), we can quickly find similar temporal context in sparse trajectories and retrieve the corresponding hidden state for prediction.
@@
Figure 3: A toy example showing the working principle of REPLAY on modeling time-varying temporal regularities of human mobility.
@@
Smoothed Timestamps further captures the temporal variation (uncertainty) of the temporal regularities. Specifically, even though the time interval in the flashback mechanism is able to capture the temporal regularity over sparse trajectories to some extent, human mobility has an intrinsic uncertainty, which implies that the regular activity of a user might be recorded at slightly different timestamps in a day in sparse trajectories, for example. We reveal in this paper that the degree of such uncertainty (or the strength of the temporal regularities) varies across time (e.g., across different time periods during a day as shown in Figure 1). Subsequently, the smoothed timestamp embeddings are used to accommodate such variations. Instead of directly using a specific timestamp embedding, we smooth it with the neighboring timestamps as shown in Figure 3, where the timestamp-specific learnable bandwidths can automatically adapt to the temporal regularities of different strengths across different timestamps.
3.5 Model Training
The training process of REPLAY follows the traditional backpropagation-through-time training process of recurrent neural networks [60]. For each prediction, we compute the cross-entropy loss between the probability distribution of POIs output by REPLAY and the ground truth POIs, which are minimized using Adam optimizer [61]. The loss function is defined as follows:
where $U$ is the number of users; $N_{j}$ is the sequence length of $j$-th user; $y_{i+1,j}$ is the one-hot vector of ground truth POI for ($i$+$1$)-th check-in of $j$-th user. $p(\hat{y}_{i+1,j}|\mathcal{H}_{i},s_{k},u_{j})$ is the predicted probability distribution given the current aggregated hidden state $\mathcal{H}_{i}$, smoothed query timestamp $s_{k}$ (transformed from time $t_{i+1}$), and user embedding $u_{j}$. For more training and implementation details of REPLAY, please refer to our code available online [^2] https://github.com/UM-Data-Intelligence-Lab/REPLAY.
@@ 4 Experiments
We conduct an extensive evaluation of REPLAY on location prediction tasks. In the following, we start by presenting our experimental setup, followed by our results and discussions.
@@ 4.1 Experiment Setup
We present our experiment setup below, including datasets, baselines, and our experiment settings.
4.1.1 Dataset
We evaluate our model on two widely used real-world datasets: Gowalla [23] and Foursquare [25], respectively. Table I shows the dataset statistics. Each mobility trace is chronologically split into 80% for training and 20% for testing.
Note that the original datasets use UTC time (Coordinated Universal Time) and it is inappropriate to directly use such time for modeling the temporal regularity of users worldwide, as the user mobility regularity depends on local time. We thus transform the UTC time of each check-in to a local time according to the time zone of the check-in location (GPS coordinates).
TABLE I: Datasets Statistics
@@ 4.1.1 Dataset
Median time between
successive check-ins
11.09 hours
(0.46 days)
16.72 hours
(0.70 days)
4.1.2 Baselines
Our baseline techniques can be divided into five types:
  • User Preference-based Methods. WRMF [62] converts implicit feedback into confidence values and employs alternating least squares for the computation of user and item latent factors. BPR proposed by [63] is a recommendation approach specifically designed for implicit feedback based on ranking, which employs a pairwise ranking loss to learn user preferences; it exploits direct user-item interaction to separate negative items from positive items to alleviate the sparsity of data and improves the performance. For these two methods, We employ the default configurations of LibRec [^3] https://github.com/guoguibing/librec.
  • Feature-based methods: Most Frequent Time (MFT) proposed by [37] based on the observation that users often interact with POIs at specific times or during particular time intervals (24 hours a day) to rank a POI. LBSN2Vec [32] employs a random walk-with-stay scheme to automatically learn embeddings from an LBSN hypergraph; for location prediction tasks, LBSN2Vec ranks POIs by maximizing the cosine similarity with user and time embeddings. In our experiments, the ratio of mobility data, the number of the window size of random walk, and the negative sample number are set to {1, 10, 10}, respectively.
  • Markov-Chain-based Methods: FPMC [50] uses matrix factorization techniques to factorize a personalized transition matrix in user-POI latent factors. The dimension of factorization and the negative sample numbers are set to {32, 10} in our experiments. PRME [49] considers spatial constraints of human mobility and extends FPMC to location prediction problems; it captures the personalized POI mobility transition patterns by learning user-POI embeddings. The component weight, the hidden state size, and the threshold are set to {0.2, 20, 360} in our experiments. TribeFlow [57] predicts user trajectories using semi-Markov transition probabilities over latent space based on a mixture model. The number of transitions, cores, batches, and iterations are set to {0.3, 20, 20, 2000}, respectively.
  • Basic RNNs: RNN [64] is a basic Recurrent Neural Network architecture; for location prediction tasks, it captures sequential patterns from check-in datasets. LSTM [4] is a variant of RNN, which contains three multiplicative gates and a memory cell to hold long-term dependencies. GRU [3] is equipped with two gates to control the information flow for the location prediction task. For all three RNN architectures and across all datasets, the hidden size is set to 10 in our experiments.
  • Spatiotemporal Sequence Models: DeepMove [58, 65] proposes an attentional recurrent network for mobility prediction from lengthy and sparse trajectories. We employ default settings as suggested by the paper, and the embedding and the hidden size are set to {50, 10}, respectively. STRNN [7] adopts time and spatial-specific transition matrices in RNN to model mobility patterns. STGN [12] adds spatial and temporal gates to the conventional LSTM to capture sequential patterns of users. STGCN [12] uses coupled input and forget gates in STGN. The hidden size is set to 10 in our experiment. STAN [51] utilizes self-attention layers along the trajectory to capture the point-to-point interaction between non-adjacent check-ins. The negative sample number, embedding size, and trajectory sequence are set to {10, 50, 100}, respectively. GetNext [66] is a graph-enhanced transformer model that exploits user global trajectory flow map for location prediction; for a fair comparison, the category of POIs are uniformly assigned to the same. The user embedding size, GCN layers’ channel numbers, POIs embedding size, and transformer encoder embedding size are set to {32, 64, 128, 1024}, respectively. Flashback [14] uses a context-aware weighting mechanism to search informative past hidden states; Graph-Flashback [17] adds knowledge graph embeddings techniques to the flashback network, where the social relationships between users are also used. We set $\alpha$ and $\beta$ as suggested by Flashback in Graph-Flashback and Flashback.
TABLE II: Overall location prediction performance, where the best-performing ones are highlighted. Methods make predictions using query time information as a query timestamp (MFT, LBSN2Vec, and REPLAY), as a query time interval (TribeFlow, DeepMove, STGN, STGCN, STAN), or as a query time threshold (PRME). Note that DeepMove predicts locations in a fixed future time slot, implicitly implying a fixed query time interval. The top two best-performing results are denoted in bold and underlined, respectively.
Method Gowalla Foursquare
Acc@1 Acc@5 Acc@10 MRR Acc@1 Acc@5 Acc@10 MRR
BPR 0.0131 0.0363 0.0539 0.0235 0.0315 0.0828 0.1143 0.0538
Feature-based Methods MFT 0.0525 0.0948 0.1052 0.0717 0.1945 0.2692 0.2788 0.2285
LBSN2Vec 0.0864 0.1186 0.1390 0.1032 0.2190 0.3955 0.4621 0.2781
PRME 0.0740 0.2146 0.2899 0.1504 0.0982 0.3167 0.4064 0.2040
TribeFlow 0.0256 0.0723 0.1143 0.0583 0.0297 0.0832 0.1239 0.0645
LSTM 0.0621 0.1637 0.2182 0.1144 0.1144 0.2949 0.3761 0.2018
GRU 0.0528 0.1416 0.1915 0.0993 0.0606 0.1797 0.2574 0.1245
DeepMove 0.0625 0.1304 0.1594 0.0982 0.2400 0.4319 0.4742 0.3270
STGN 0.0624 0.1586 0.2104 0.1125 0.2094 0.4734 0.5470 0.3283
STGCN 0.0546 0.1440 0.1932 0.1017 0.1878 0.4502 0.5329 0.3062
STAN 0.0891 0.2096 0.2763 0.1523 0.2265 0.4515 0.5310 0.3420
GETNext 0.0912 0.2003 0.2487 0.1484 0.1862 0.4702 0.5763 0.3153
Flashback 0.1158 0.2754 0.3479 0.1925 0.2496 0.5399 0.6236 0.3805
REPLAY 0.1866 0.3476 0.4124 0.2635 0.3529 0.5953 0.6648 0.4638
4.1.3 Experiment Settings
We developed our model using the PyTorch framework and conducted experiments on the following hardware platform (CPU: Intel(R) Xeon(R) Gold 5320, GPU: NVIDIA GeForce RTX 3090). We evaluate REPLAY and the baselines in the location prediction task, aiming to predict the next POI a user will visit, based on her historical check-in sequence and a query time, as illustrated in Figure 2. We use two widely used evaluation metrics: Mean Reciprocal Rank (MRR) and average Accuracy@N (Acc@N), we adopt N = {1, 5, 10} in our experiment. We empirically set the dimension of hidden states and all (POI, timestamp, and user) embedding sizes as 10, and the temporal decay factor $\alpha$ and spatial decay factor $\beta$ following the suggested settings in [14], i.e., $\alpha=0.1$ on both datasets, $\beta=1000$ and $100$ on Gowalla and Foursquare datasets, respectively, and the distance to flashback as 20. We implement our REPLAY with different RNN architectures (including vanilla RNN, GRU and LSTM), and systematically compare their performance in an ablation study below; in other experiments, we report the results of REPLAY using the best RNN architecture setting (vanilla RNN), if not specified otherwise.